# The quantum Newton’s bucket: Active and passive rotations in quantum theory

### v1

@@
Augusto Facundes da Silva, Kayman Jhosef Goncalves, Giorgio Torrieri
@@
I Introduction
Isaac Newton, in his work “Philosophiæ Naturalis Principia Mathematica” Newton (1687); Assis (1997), discusses a thought experiment in which a cylindrical container, such as a bucket, is partially filled with a fluid, like water, and is suspended in the air by a rope. After twisting the rope sufficiently and releasing it, the bucket rotates with a constant angular velocity. At the beginning of the motion, only the bucket rotates, while the water remains at rest. Then, the water gradually starts rotating with the same angular velocity as the bucket, forming a surface in the liquid that takes the shape of a paraboloid of revolution—different from when the water is at rest, where the surface is flat. This curvature remains for some time after the bucket stops rotating, returning to a flat surface when the water also comes to rest. The thought experiment demonstrates that the physical effects of rotation are absolute, not relative. Examining each stage in detail:
@@ I Introduction
This is quite different from the “relativity of velocity” in Galileo’s principle of relativity and justifies Newton’s claim that space is, in some sense, absolute. Mach attempted to define “absolute acceleration” in terms of the asymptotic boundary (“distant stars”), whereas Einstein postulated that acceleration is defined in terms of local spacetime, dynamically determined by gravity. The relationship between the two definitions, and the implications for dynamics, gravity and cosmology are still subject to debate in the literature Rothman (2017).
We were reminded of these classic conudrums by the recent interest in the literature of “QCD in rotating referece frames” Braguta et al. (2023); Chernodub (2021); Ayala et al. (2021); Prokhorov et al. (2020); Yamamoto and Hirono (2013); Aristova et al. (2016); Wang et al. (2019) motivated by the experimental finding of vortical effects in Quark-Gluon plasma Becattini et al. (2021a); Niida and Voloshin (2024); Wang (2017). A variety of both perturbative and non-perturbative approaches have been used to calculate quantities in a rotating medium. Yet the way rotation was imposed was essentially to work with a metric put in by hand, which corresponds to a passive rotation. Of course, in a laboratory systems rotate because there is something in their dynamics that makes them rotate. To what extent are the two equivalent? In the rest of this work we shall investigate this “elementary” but surprisingly subtle question in a variety of systems, and will try to establish criteria for this equivalence.
One has to be clear as to what “active” and “passive” mean mathematically. An active rotation is one where the rotation is due to the Hamiltonian of the system, so
$H\rightarrow H+H_{\text{rotation}}\phantom{AA},\phantom{AA}|\psi\rangle\rightarrow\exp[iH_{\text{rotation}}t]\ket{\psi}$ (1)
or equivalently, given a complete set of Eigenstates |$\ket{\psi}_{i}$ and coefficients normalized $c_{ij}$
$\hat{\rho}=c_{ij}\ket{\psi_{i}}\bra{\psi_{j}}\phantom{AA},\phantom{AA}\frac{d\hat{\rho}}{dt}=i[\hat{\rho},\hat{H}]$ (2)
where the classical counterpart of the hamiltonian $\hat{H}$ includes a rotation potential and the state of the system is close to ”a wavepacket”.
In a passive rotation the system is stationary, and the detector rotates around the system. This can be accomplished by
@@
$H\rightarrow H_{0}-\bm{\omega}\cdot\mathbf{J}\phantom{AA},\phantom{AA}[\mathbf{J},H_{0}]=0,$ (4)
both rotations are the same, because the “passive” rotation matrix is exactly the same as the difference between the Hamiltonian evolutions (“before” and “after” the shift). A crucial requirement, quintessentially quantum, is however $[\mathbf{J},H_{0}]=0$ as we shall see.
@@
II Quantum mechanics
In the following, We consider a ”high-lying” energy state where one can think of ”a rotating wavepacket” rather than a wavefunction, under a variety of potentials. Physically, this system is close to what is known as a Rydberg atom Gallagher (1994)
II.1 Coulombic potential
@@ II.1.1 Active rotation
where $|r\rangle=|n,\ell,m_{\ell}\rangle$ represents the states in the position space, $|S\rangle=|s,m_{s}\rangle$, $s=\{1,2\}$, the states in the spin space and $\rho_{m_{1},m_{2}}$ the normalization coefficients of the density matrix in $t=0$. Rewriting the tensor product in (6) in the form
@@
$\boxed{\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i(H_{M_{1}}-H_{M_{2}})\frac{t}{\hbar}}}.$ (13)
The energy eigenvalues $H_{M_{i}}$ can be determined for a particle under rotation in a cylindrical box with radius $R\gg\varrho$, which corresponds, with the potential used here, to an hydrogen atom under rotation such that it satisfies $|n,\ell,m_{\ell}\rangle=|n,\ell,m_{\ell}\rangle_{H}=|n,\ell,m_{\ell};t=0\rangle_{S}=\psi(\varrho,\phi,z)$, which is given by
$\psi(\varrho,\phi,z)=A_{n,\ell,m}\left(\frac{\varrho}{na}\right)^{m}\left(\frac{z}{na}\right)^{\ell-m}e^{-\frac{\sqrt{\varrho^{2}+z^{2}}}{na}}e^{im\phi},$ (14)
where $A_{n,\ell,m}$ is the normalization constant and $a$ the Bohr radius, whose energy eigenvalues are given by
$H_{M_{i}}=-\frac{\alpha m}{2\hbar^{2}n^{2}}-\hbar\omega_{i}M_{i}.$ (15)
@@
$\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i\left(-\frac{\alpha m}{2\hbar^{2}n^{2}}-\hbar\omega_{1}M_{1}+\frac{\alpha m}{2\hbar^{2}n^{2}}+\hbar\omega_{2}M_{2}\right)\frac{t}{\hbar}}$
@@
$\boxed{\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i\left(M_{2}-M_{1}\right)\omega_{z}t}}.$ (16)
II.1.2 Passive rotation
In the reference frame of the particle, the detector rotates around the $z$-axis by and angle $\phi$ (via the action of the rotation operator $U(\phi,t)=D_{z}(\phi,t)=e^{-iJ_{z}\frac{\phi}{t}}$, where $J_{z}$ is the angular momentum operator in the $z$-direction). Following the equation Eq. (1),
@@ II.1.2 Passive rotation
$\boxed{\displaystyle\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i(J_{z,M_{1}}\omega_{z_{1}}-J_{z,M_{2}}\omega_{z_{2}})\frac{t}{\hbar}}.}$ (19)
@@
$\boxed{\displaystyle\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i(M_{1}-M_{2})\omega_{z}t}.}$ (20)
Comparing the equations Eq. (16) and Eq. (20), the density matrix $\hat{\rho}(t)$ has the same form, Except for the sign corresponding to the difference $M_{2}-M_{1}=-(M_{1}-M_{2})$ in the exponential term, which arises because the rotation is viewed from different reference frames. This indicates that the active and passive rotations of a generic spin particle under the action of a Coulomb potential are equivalent.
II.2 Position-varying magnetic field in two dimensions
II.2.1 Active rotation
In the presence of a position-dependent magnetic potential $\bm{B}(\bm{r})$, the Hamiltonian of the system under active rotation acquires the contribution
@@ II.2.1 Active rotation
where $\bm{B(r)}=\displaystyle\left(\Omega_{1}+\frac{\Omega_{2}}{r}+\frac{\Omega_{3}}{r^{2}}\right)\hat{\bm{z}}$ e $\hat{\bm{\mu}}=\gamma\frac{q}{2m}\hat{\bm{S}}$ is the magnetic moment operator, with $\gamma$ being the gyromagnetic ratio and $q$ the particle’s charge.
To find the energy eigenvalues of this system, the Hamiltonian is applied Eq. (21) in the Schrödinger equation, resulting in
@@
$\frac{1}{2mr^{2}}\frac{d}{dr}\left(r^{2}\frac{dR(r)}{dr}\right)-\left[-\frac{\alpha}{r}-\omega_{z}M_{j}-\gamma\frac{q}{2m}M_{j}\left(\Omega_{1}+\frac{\Omega_{2}}{r}+\frac{\Omega_{3}}{r^{2}}\right)-E\right]R(r)=0$ (24)
@@
$\frac{d^{2}\chi(x)}{dx^{2}}+\frac{2x}{x^{2}}\frac{d\chi(x)}{dx}+\frac{2m}{x^{4}}\bigg{[}-\omega_{z}M_{j}-\gamma\frac{q}{2m}M_{j}\Omega_{1}-E-\left(\alpha+\gamma\frac{q}{2m}M_{j}\Omega_{2}\right)x-\\ \gamma\frac{q}{2m}M_{j}\Omega_{3}x^{2}\bigg{]}\chi(x)=0,$ (25)
@@
$H_{0}=\omega_{z}M_{j}+\gamma\frac{q}{2m}M_{j}\Omega_{1}+E\phantom{AA},\phantom{AA}H_{1}=-\alpha-\gamma\frac{q}{2m}M_{j}\Omega_{2}\phantom{AA},\phantom{AA}H_{2}=-\gamma\frac{q}{2m}M_{j}\Omega_{3}.$ (27)
By the Nikiforov-Uvarov method De Moura et al. (2023a), the energy eigenvalues will be given by
$\sqrt{H_{0}}=\frac{H_{1}}{(1+2n)\pm\sqrt{1-4H_{2}}},$ (28)
@@
$\omega_{z}M_{j}+\gamma\frac{q}{2m}M_{j}\Omega_{1}+E=\left(\frac{-\alpha-\gamma\frac{q}{2m}M_{j}\Omega_{2}}{(1+2n)+\sqrt{1+2\gamma\frac{q}{m}M_{j}}}\right)^{2}$ (29)
$E=\left(\frac{-\alpha-\gamma\frac{q}{2m}M_{j}\Omega_{2}}{(1+2n)+\sqrt{1+2\gamma\frac{q}{m}M_{j}}}\right)^{2}-M_{j}\left(\omega_{z}+\gamma\frac{q}{2m}\right)$ (30)
Thus, the density matrix operator given by equation Eq. (13) for the active rotation will be given by
$\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}\times\\ \exp\Bigg{\{}i\Bigg{[}\left(\frac{-\alpha-\gamma\frac{q}{2m}M_{1}\Omega_{2}}{(1+2n)+\sqrt{1+2\gamma\frac{q}{m}M_{1}}}\right)^{2}-\left(\frac{-\alpha-\gamma\frac{q}{2m}M_{2}\Omega_{2}}{(1+2n)+\sqrt{1+2\gamma\frac{q}{m}M_{2}}}\right)^{2}+\\ M_{2}\left(\omega_{z}+\gamma\frac{q}{2m}\right)-M_{1}\left(\omega_{z}+\gamma\frac{q}{2m}\right)\Bigg{]}t\Bigg{\}}.$ (31)
For this system, the states correspond to the wave function solution of the time-independent Schrödinger equation associated with equation Eq. (22), which can be written as
@@
II.2.2 Passive rotation
In the system under passive rotation, the density operator takes the same form of equation Eq. (20), derived in an equivalent way.
Thus, for the rotations to be equivalent, the exponential term in equation Eq. (31) must be equal to $(M_{1}-M_{2})\tilde{\omega}_{z}$. The issue is that, for this to occur in the passive rotation for this configuration, there must be an angular velocity $\tilde{\omega}_{z}$ that depends on the quantum numbers $n$ and $m_{s}$, which does not seem possible. Therefore, for a space-varying magnetic field in the form defined here, the active and passive rotations of a generic spin particle are not equivalent.
II.3 Rotating cylindrical wells
Now, consider the case of a non-relativistic cylindrical well of radius $R$ and constant depth $U_{0}$, rotating with constant angular velocity $\omega$ about the $z$-axis, passing through its center of mass, as discussed in Buzzegoli and Tuchin (2023). The Hamiltonian of this system reads

### v2

@@
Augusto Facundes, Kayman J. Gonçalves, Giorgio Torrieri
@@
1 Introduction
Isaac Newton, in his work “Philosophiæ Naturalis Principia Mathematica” [Newton(1687), Assis(1997)], discusses a thought experiment in which a cylindrical container, such as a bucket, is partially filled with a fluid, like water, and is suspended in the air by a rope. After twisting the rope sufficiently and releasing it, the bucket rotates with a constant angular velocity. At the beginning of the motion, only the bucket rotates, while the water remains at rest. Then, the water gradually starts rotating with the same angular velocity as the bucket, forming a surface in the liquid that takes the shape of a paraboloid of revolution—different from when the water is at rest, where the surface is flat. This curvature remains for some time after the bucket stops rotating, returning to a flat surface when the water also comes to rest. The thought experiment demonstrates that the physical effects of rotation are absolute, not relative. Examining each stage in detail:
@@ 1 Introduction
This is quite different from the “relativity of velocity” in Galileo’s principle of relativity and justifies Newton’s claim that space is, in some sense, absolute. Mach attempted to define “absolute acceleration” in terms of the asymptotic boundary (“distant stars”), whereas Einstein postulated that acceleration is defined in terms of local spacetime, dynamically determined by gravity. The relationship between the two definitions, and the implications for dynamics, gravity and cosmology are still subject to debate in the literature [Rothman(2017)].
We were reminded of these classic conudrums by the recent interest in the literature of “QCD in rotating referece frames” [Braguta et al.(2023)Braguta, Kudrov, Roenko, Sychev, and Chernodub, Chernodub(2021), Ayala et al.(2021)Ayala, Hernández, Raya, and Zamora, Prokhorov et al.(2020)Prokhorov, Teryaev, and Zakharov, Yamamoto and Hirono(2013), Aristova et al.(2016)Aristova, Frenklakh, Gorsky, and Kharzeev, Wang et al.(2019)Wang, Wei, Li, and Huang] motivated by the experimental finding of vortical effects in Quark-Gluon plasma [Becattini et al.(2021a)Becattini, Liao, and Lisa, Niida and Voloshin(2024), Wang(2017)]. A variety of both perturbative and non-perturbative approaches have been used to calculate quantities in a rotating medium. Yet the way rotation was imposed was essentially to work with a metric put in by hand, which corresponds to a passive rotation. Of course, in a laboratory, systems rotate because there is something in their dynamics that makes them rotate. To what extent are the two equivalent? Some of the works in [Braguta et al.(2023)Braguta, Kudrov, Roenko, Sychev, and Chernodub, Chernodub(2021), Ayala et al.(2021)Ayala, Hernández, Raya, and Zamora, Prokhorov et al.(2020)Prokhorov, Teryaev, and Zakharov, Yamamoto and Hirono(2013), Aristova et al.(2016)Aristova, Frenklakh, Gorsky, and Kharzeev, Wang et al.(2019)Wang, Wei, Li, and Huang] use the “rigid body“ approximation (where, for conserved angular momentum, active and passive rotations are indeed trivially equivalent), but since rigidity is violated both by relativity and quantum mechanics, there is no clear sense this is a well-defined effective theory.
In the rest of this work we shall investigate this “elementary”, but surprisingly subtle question in a variety of systems, and will try to establish criteria for this equivalence.
One has to be clear as to what “active” and “passive” mean mathematically, especially as the concept of accelleration in general only makes sense in the semi-classical limit (the Newton relation $d\hat{p}/dt=mdV/d\hat{x}$ becomes ill-defined if commutators between position and momentum are non-negligible).
An active rotation is one where the rotation is due to the Hamiltonian of the system, so
[…]
where the classical counterpart of the Hamiltonian $\hat{H}$ includes a rotation potential and the state of the system is close to a “wavepacket”.
In a passive rotation, the system is stationary, and the detector rotates around the system. This can be accomplished by
@@
$\hat{H}\rightarrow\hat{H}_{0}-\bm{\omega}\cdot\mathbf{\hat{J}}\phantom{AA},\phantom{AA}[\mathbf{\hat{J}},\hat{H}_{0}]=0,$ (4)
both rotations are the same, because the “passive” rotation matrix is exactly the same as the difference between the Hamiltonian evolutions (“before” and “after” the shift). As we shall see, a crucial requirement is that $\mathbf{J},H_{0}]=0$; This requirement is “quintessentially“ quantum, as it reflect the relationship between lack of commutation and experimental uncertainity.
@@
2 Quantum mechanics
In the following, we consider a “high-lying” energy state where one can think of “a rotating wavepacket” rather than a wavefunction, under a variety of potentials. Physically, this system is close to what is known as a Rydberg atom [Gallagher(1994)]
2.1 Coulombic potential
@@ 2.1.1 Active rotation
where $|r\rangle=|n,\ell,m_{\ell}\rangle$ represents the states in the position space, $|S\rangle=|s,m_{s}\rangle$ the states in the spin space and $\rho_{m_{1},m_{2}}$ the normalization coefficients of the density matrix at $t=0$. Rewriting the tensor product in (6) in the form
@@
$\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i(H_{M_{1}}-H_{M_{2}})\frac{t}{\hbar}}.$ (13)
The energy eigenvalues $H_{M_{i}}$ can be determined for a particle under rotation in a spherical box, which corresponds, with the potential used here, to a hydrogen-like atom under rotation such that it satisfies $|n,\ell,m_{\ell}\rangle=|n,\ell,m_{\ell}\rangle_{H}=|n,\ell,m_{\ell};t=0\rangle_{S}=\psi(r,\theta,\phi)$, which is given by
$\psi(r,\theta,\phi)=A_{n,\ell,m}e^{-r/na_{\kappa}}\left(\frac{2r}{na_{\kappa}}\right)^{\ell}L^{2\ell+1}_{n-\ell-1}\left(\frac{2r}{na_{\kappa}}\right)Y^{m}_{\ell}(\theta,\phi),$ (14)
where $A_{n,\ell,m}$ is the normalization constant and $a_{\kappa}$ the generalized Bohr radius, such that $\kappa\equiv|Q_{p}Q_{N}|$, with $Q_{p}$ and $Q_{N}$ the charges of the generalized particle and generalized nucleus, respectively. The energy eigenvalues are then given by
$H_{M_{i}}=-\frac{\alpha^{2}m}{2\hbar^{2}n^{2}}-\hbar\omega_{i}M_{i}.$ (15)
@@
$\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i\left(-\frac{\alpha^{2}m}{2\hbar^{2}n^{2}}-\hbar\omega_{1}M_{1}+\frac{\alpha^{2}m}{2\hbar^{2}n^{2}}+\hbar\omega_{2}M_{2}\right)\frac{t}{\hbar}}$
@@
$\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i\left(M_{2}-M_{1}\right)\omega_{z}t}.$ (16)
2.1.2 Passive rotation
In the reference frame of the particle, the detector rotates around the $z$-axis by an angle $\phi$ (via the action of the rotation operator $U(\phi,t)=D_{z}(\phi,t)=e^{-iJ_{z}\frac{\phi}{t}}$, where $J_{z}$ is the angular momentum operator in the $z$-direction). Following Eq. (1),
@@ 2.1.2 Passive rotation
$\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i(J_{z,M_{1}}\omega_{z_{1}}-J_{z,M_{2}}\omega_{z_{2}})\frac{t}{\hbar}}.$ (19)
@@
$\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}e^{i(M_{1}-M_{2})\omega_{z}t}.$ (20)
Comparing Eq. (16) and Eq. (20), the density matrix $\hat{\rho}(t)$ has the same form, except for the sign corresponding to the difference $M_{2}-M_{1}=-(M_{1}-M_{2})$ in the exponential term, which arises because the rotation is viewed from different reference frames. This indicates that the active and passive rotations of a generic spin particle under the action of a Coulomb potential are equivalent.
2.2 Position-varying magnetic field in two dimensions
2.2.1 Active rotation
In the presence of a position-dependent magnetic field $\bm{B}(\bm{r})$, the Hamiltonian of the system under active rotation acquires the contribution
@@ 2.2.1 Active rotation
where $\bm{B(r)}=\displaystyle\left(\Omega_{1}+\frac{\Omega_{2}}{r}+\frac{\Omega_{3}}{r^{2}}\right)\hat{\bm{z}}$ and $\hat{\bm{\mu}}=\gamma\frac{q}{2m}\hat{\bm{S}}$ is the magnetic moment operator, with $\gamma$ being the gyromagnetic ratio and $q$ the particle’s charge.
To find the energy eigenvalues of this system, the Hamiltonian is applied in the Schrödinger equation, resulting in
@@
$\frac{1}{r^{2}}\frac{d}{dr}\left(r^{2}\frac{dR(r)}{dr}\right)-{\color[rgb]{0,0,0}2m}\left[{\color[rgb]{0,0,0}\frac{\ell(\ell+1)}{r^{2}}}-\frac{\alpha}{r}-\omega_{z}M_{j}-\gamma\frac{q}{2m}M_{j}\left(\Omega_{1}+\frac{\Omega_{2}}{r}+\frac{\Omega_{3}}{r^{2}}\right)-E\right]R(r)=0$ (24)
@@
$\frac{d^{2}\chi(x)}{dx^{2}}+\frac{2x}{x^{2}}\frac{d\chi(x)}{dx}+\frac{2m}{x^{4}}\bigg\{\omega_{z}M_{j}+\gamma\frac{q}{2m}M_{j}\Omega_{1}+E+\left(\alpha+\gamma\frac{q}{2m}M_{j}\Omega_{2}\right)x+\\ \left[\gamma\frac{q}{2m}M_{j}\Omega_{3}-{\color[rgb]{0,0,0}\frac{\ell(\ell+1)}{2m}}\right]x^{2}\bigg\}\chi(x)=0,$ (25)
@@
$H_{0}={\color[rgb]{0,0,0}-2m}\left(\omega_{z}M_{j}+\gamma\frac{q}{2m}M_{j}\Omega_{1}+E\right)\phantom{AA},\phantom{AA}H_{1}={\color[rgb]{0,0,0}2m}\left(\alpha+\gamma\frac{q}{2m}M_{j}\Omega_{2}\right)\phantom{AA},\phantom{AA}\\ H_{2}={\color[rgb]{0,0,0}2m}\left[\gamma\frac{q}{2m}M_{j}\Omega_{3}-{\color[rgb]{0,0,0}\frac{\ell(\ell+1)}{2m}}\right].$ (27)
By the Nikiforov-Uvarov method [De Moura et al.(2023a)De Moura, Gonçalves, and Torrieri, Yazdankish(2020)], the energy eigenvalues will be given by
$\sqrt{H_{0}}=\frac{H_{1}}{1+2n\pm\sqrt{1-4H_{2}}},$ (28)
@@
${\color[rgb]{0,0,0}2m}\left(\omega_{z}M_{j}+\gamma\frac{q}{2m}M_{j}\Omega_{1}+E\right)=\left[\frac{{\color[rgb]{0,0,0}2m}\left(\alpha+\gamma\frac{q}{2m}M_{j}\Omega_{2}\right)}{1+2n+\sqrt{1-{\color[rgb]{0,0,0}2m}\left(2\gamma\frac{q}{m}M_{j}\Omega_{3}-2{\color[rgb]{0,0,0}\frac{\ell(\ell+1)}{m}}\right)}}\right]^{2}$ (29)
$E=2m\left[\frac{\alpha+\gamma\frac{q}{2m}M_{j}\Omega_{2}}{1+2n+\sqrt{1-4\left(\gamma qM_{j}\Omega_{3}-{\color[rgb]{0,0,0}\ell(\ell+1)}\right)}}\right]^{2}-M_{j}\left(\omega_{z}+\gamma\frac{q}{2m}\Omega_{1}\right)$ (30)
Thus, the density matrix operator given by Eq. (13) for the active rotation will be given by
$\hat{\rho}(t)=\sum_{m_{\ell},m_{s}}C^{J,M_{1}}_{s,m_{\ell},\ell,m_{1}}C^{J,M_{2}}_{s,m_{\ell},\ell,m_{2}}\rho_{m_{1},m_{2}}\exp\Bigg\{i\Bigg[2m\left(\frac{\alpha+\gamma\frac{q}{2m}M_{1}\Omega_{2}}{1+2n+\sqrt{1-4(\gamma qM_{1}\Omega_{3}-\ell(\ell+1))}}\right)^{2}-\\ 2m\left(\frac{\alpha+\gamma\frac{q}{2m}M_{2}\Omega_{2}}{1+2n+\sqrt{1-4(\gamma qM_{2}\Omega_{3}-\ell(\ell+1))}}\right)^{2}+\left(M_{2}-M_{1}\right)\left(\omega_{z}+\gamma\frac{q}{2m}\Omega_{1}\right)\Bigg]t\Bigg\}.$ (31)
For this system, the states correspond to the wave function solution of the time-independent Schrödinger equation associated with Eq. (22), which can be written as
@@
2.2.2 Passive rotation
In the system under passive rotation, the density operator takes the same form of Eq. (20), derived in an equivalent way. Thus, for the rotations to be equivalent, the exponential term in Eq. (31) must be equal to $(M_{1}-M_{2})\tilde{\omega}_{z}$. The issue is that, for this to occur in the passive rotation for this configuration, there must be an angular velocity $\tilde{\omega}_{z}$ that depends on the quantum numbers $n$ and $m_{s}$, which is not possible since, in the rotating frame, $m_{s}$ is not a good quantum number. Therefore, for a space-varying magnetic field in the form defined here, the active and passive rotations of a generic spin particle are not equivalent.
2.3 Rotating cylindrical wells
Now, consider the case of a non-relativistic cylindrical well of radius $R$ and constant depth $U_{0}$, rotating with constant angular velocity $\omega$ about the $z$-axis, passing through its center of mass, as discussed in [Buzzegoli and Tuchin(2023)]. The Hamiltonian of this system reads